### v1

@@ Abstract
…..We study a compositional variant of kernel ridge regression in which the predictor is applied to a coordinate-wise reweighting of the inputs. Formulated as a variational problem, this model provides a simple testbed for feature learning in compositional architectures. From the perspective of variable selection, we show how relevant variables are recovered while noise variables are eliminated. We establish guarantees showing that both global minimizers and stationary points discard noise coordinates when the noise variables are Gaussian distributed. A central finding is that $\ell_{1}$-type kernels, such as the Laplace kernel, succeed in recovering features contributing to nonlinear effects at stationary points, whereas Gaussian kernels recover only linear ones.
@@ 1 Introduction
Our starting point is the classical kernel ridge regression (KRR) framework, which aims to learn a nonlinear function $f$ that predicts the response $Y\in\mathbb{R}$ from the covariate $X\in\mathbb{R}^{d}$, where $(X,Y)\sim\mathbb{P}$ [2]. The function $f$ is chosen from a reproducing kernel Hilbert space $H$ (e.g., Sobelev space). In KRR, the learning problem is formulated as the solution to the variational problem
@@ 1.2 Related Work
    Another recent line of work centers on the so-called AGOP (Average Gradient Outer Product) method [12]. This line of work also adopts the parameterization $f(\beta\circ X)$ (and $f(UX)$) within kernel ridge regression and can be viewed as operating on the same objective $\mathcal{J}$. AGOP relies on an iterative update scheme $\beta^{(t+1)}={\rm AGOP}(\beta^{(t)})$ (and $U^{(t+1)}={\rm AGOP}(U^{(t)})$ resp.) designed to extract a low-dimensional set of relevant features, and has been studied both theoretically—with generalization error bounds that outperform that of standard kernel methods—and empirically, with remarkable success on tasks such as tabular dataset predictions [12, 16]. Our study takes a complementary perspective. Instead of analyzing a specific iterative scheme, we focus on the optimization landscape of $\mathcal{J}$ itself and establish population-level guarantees on feature recovery for its global minimizers and stationary points. These results hold under more relaxed distributional assumptions and highlight how the geometry of $\mathcal{J}$ shapes variable selection independently of any algorithmic choices. A central insight of this analysis is that the choice of kernel is crucial: whereas existing guarantees primary focus on smooth kernels, we show that nonsmooth $\ell_{1}$-type kernels such as the Laplace kernel enables the recovery of nonlinear main effects that Gaussian kernels miss. Interestingly, a recent algorithmic development on AGOP—namely xRFM [17], a state-of-the-art method for tabular datasets—demonstrates empirically that adopting the Laplace kernel improves generalization over other kernels (e.g., Gaussian), aligning with the theoretical insights developed in this paper.
@@ 1.4 Notation
Let $x\in\mathbb{R}^{d}$ be a vector. We write $x=(x_{1},\ldots,x_{d})$, where $x_{i}$ denotes the $i$-th coordinate of $x$, and define its support as $\mathop{\rm supp}(x):=\{i\in\{1,\ldots,d\}:x_{i}\neq 0\}$. For a subset $S\subset\{1,\ldots,d\}$, we write $x_{S}\in\mathbb{R}^{|S|}$ for the subvector of $x$ containing coordinates $\{x_{i}:i\in S\}$. We use $\|\cdot\|_{p}$ to denote the $\ell_{p}$ norm on $\mathbb{R}^{d}$. We use $e_{i}\in\mathbb{R}^{d}$ to denote the $i$-th standard basis vector. We use $\bar{z}$ to denote the complex conjugate of a complex number $z\in\mathbb{C}$.
@@ 2.1 Translation Invariant RKHS
Let $H$ be a Hilbert space of real-valued continuous function on $\mathbb{R}^{d}$ associated to a translation-invariant kernel, and $C(\mathbb{R}^{d})$ denote the space of continuous functions equipped with $L_{\infty}$ norm. The space $H$ is defined in terms of the Fourier transform by
$H:=\left\{f:f\in C(\mathbb{R}^{d})\cap L_{2}(\mathbb{R}^{d}),\left\|{f}\right\|_{H}^{2}=\int_{\mathbb{R}^{d}}\frac{|\hat{f}|^{2}(\omega)}{k(\omega)}d\omega<\infty\right\}$
@@
Lemma 1 (Continuous embedding).
For any $f\in H$, $\left\|{f}\right\|_{L_{\infty}}\leq\left\|{f}\right\|_{H}$, and $\lim_{x\to\infty}f(x)=0$.
@@
By the Riemann-Lebesgue lemma, we get $\lim_{x\to\infty}f(x)=0$. ∎
@@
Suppose this is not true. There is nonzero $f\in L_{2}(\mu)$ such that $\int fhd\mu=0$ for all $h\in H$. In particular, taking $h(\cdot)=K(\cdot-z)\in H$, we obtain $\int f(x)K(x-z)d\mu(x)=0$ for all $z\in\mathbb{R}^{d}$. Using the Fourier inversion formula, $K(x-z)=\int k(\omega)e^{2\pi i\langle x-z,\omega\rangle}d\omega$, we get
$\begin{split}\iint f(x)k(\omega)e^{2\pi i\langle x-z,\omega\rangle}d\omega d\mu(x)=0~~~\forall z\in\mathbb{R}^{d}.\end{split}$
Define the Fourier transform of $f$ under $\mu$ as $F(\omega):=\int f(x)e^{2\pi i\langle x,\omega\rangle}d\mu(x)$. Then the above equation becomes
$\int F(\omega)k(\omega)e^{-2\pi i\langle z,\omega\rangle}d\omega=0~~~\forall z\in\mathbb{R}^{d}.$
That is, the inverse Fourier transform of $F(\omega)k(\omega)$ is identically zero. Since $k(\omega)>0$, this implies $F(\omega)=0$ almost everywhere $\omega$, so $f=0$ in $L_{2}(\mu)$—a contradiction. ∎
@@ 2.3 Continuity in Parameter
A fundamental property is that the minimizer $f_{\beta}$ and the objective $\mathcal{J}$ depend continuously on the parameter $\beta$. This fact essentially follows from the reinterpretation of the $f(UX)$ model in related work [14]; for completeness, we include a self-contained proof in this manuscript.
@@
Since $f_{\beta,\lambda}\in H$ is continuous and uniformly bounded in $\mathbb{R}^{d}$, we get $\mathcal{I}(f_{\beta,\lambda},\beta^{\prime},\lambda)\to\mathcal{I}(f_{\beta,\lambda},\beta,\lambda)=\mathcal{J}(\beta,\lambda)$ as $\beta^{\prime}\to\beta$. Also, by Lemma 8, we have $\mathcal{J}(\beta^{\prime},\lambda)\to\mathcal{J}(\beta,\lambda)$. This implies
@@ 2.4 First Variation
Related formulas have appeared in the author’s earlier work on $f(UX)$ model [14], though under stronger smoothness assumptions, i.e., full Fréchet differentiability on the map $\beta\mapsto K_{\beta}(x,x^{\prime})$ for fixed $x,x^{\prime}$, and in more specialized settings where $K$ is assumed to be a radial basis kernel.
@@ 2.4.1 Examples
Theorem 1 and the above examples illustrate that, in many scenarios of interest, the directional derivative $\mathrm{D}\mathcal{J}(\beta,\lambda)[v]$ exists for all $\beta\in\mathbb{R}^{d}$ and all directions $v\in\mathbb{R}^{d}$. Motivated by these results, throughout the paper we will be only interested in scenarios where $\mathrm{D}\mathcal{J}(\beta,\lambda)[v]$ exists for all $\beta$ and directions $v$.
@@ 3.1 Existence
Continuity of $X$ serves as a sufficient condition for $\mathcal{J}$ to admit a finite minimizer. In the presence of discrete variables, it may happen that $\limsup_{\beta\to\infty}\mathcal{J}(\beta,\lambda)<\mathbb{E}[Y^{2}]$ (see [14, Section 5]). However, we are not aware of any situation where there is a global minimizer at infinity—a question we leave open.
@@ 3.2.1 Core Sufficient Feature Set
To enable a rigorous study of feature learning, we need to focus on settings where a well-defined target structure exists. Therefore, we assume the existence of core sufficient feature set throughout:
@@
Given Assumption 3.1, there is a meaningful statistical target—the core sufficient feature set $S_{*}$, which captures the minimal collection of input features required to predict the response $Y$. The problem of variable selection concerns whether minimizing $\mathcal{J}$ successfully recovers $S_{*}$. Understanding when and why this recovery occurs is central to unraveling how compositional model $f(\beta\circ X)$ learns structure from data.
@@ 3.2.2 Irrelevant Feature Elimination
We first study one aspect of feature learning: when does the global minimizer $\beta_{*}$ eliminate irrelevant features by setting coordinates outside the core sufficient feature set $S_{*}$ to zero? Such sparsity shows that the model can suppress irrelevant inputs automatically.
@@
This assumption states that the relevant features in $X$ for predicting YY—namely, those in the core sufficient feature set $S_{*}$—are independent of the remaining features $X_{S_{*}^{c}}$. Such structural independence avoids potential statistical confounding, making it clear in principle to distinguish relevant features based solely on their predictive contribution.
@@
Assume $\mathbb{E}[Y|X]\neq 0$. Let $\mathbf{1}\in\mathbb{R}^{d}$ denote the all-ones vector. We claim that $\mathcal{J}(\mathbf{1},\lambda)<\mathbb{E}[Y^{2}]$. Suppose instead that $\mathcal{J}(\mathbf{1},\lambda)=\mathbb{E}[Y^{2}]$. Then the minimizer $f_{\mathbf{1},\lambda}$ must be identically zero, which by Lemma 4 implies $\mathbb{E}[Yg(X)]=0$ for all $g\in H$. Since $H$ is dense in $L_{2}(\mathbb{P}_{X})$ by Lemma 2, it follows that $\mathbb{E}[Yg(X)]=0$ for all $g\in L_{2}(\mathbb{P}_{X})$, and hence $\mathbb{E}[Y|X]=0$ almost surely—contradicting our assumption $\mathbb{E}[Y|X]\neq 0$. We conclude that $\mathcal{J}(\mathbf{1},\lambda)<\mathbb{E}[Y^{2}]$, and thus $\inf_{\beta\in\mathbb{R}^{d}}\mathcal{J}(\beta,\lambda)<\mathbb{E}[Y^{2}]$.
@@ 3.2.3 Relevant Feature Recovery
On the other hand, since $H$ is dense in $L_{2}(\mathbb{P}_{X})$ by Lemma 2, where $\mathbb{P}_{X}$ denotes the distribution of $X$, the regression function $x\mapsto\mathbb{E}[Y|X=x]$ can be approximated arbitrarily well in $\mathcal{L}_{2}(\mathbb{P}_{X})$ by functions in $H$. In particular, there exists $f_{\varepsilon}\in H$ such that
@@ 4 Stationary Points and Irrelevant Feature Elimination
Unlike global minimizers, stationary points are more difficult to characterize, and analyzing them requires more delicate analytical techniques. To proceed, we divide the analysis into two parts. This section focuses on the following question:
@@
Fix $\beta\in\mathbb{R}^{d}$. Since the problem is translation-invariant, we can without loss of generality assume $\mathbb{E}[X]=0$. Let $\xi\in\mathbb{R}^{d}$ be a Gaussian random vector independent of $(X,Y)$ with
@@
Let $\pi_{s}$ denote the probability measure of the random vector $\sqrt{s}\,\beta_{S_{*}^{c}}\circ X_{S_{*}^{c}}\in\mathbb{R}^{|S_{*}^{c}|}$. For every function $f\in H$ and $s>0$, we define $P_{s}$ as the convolution type operator:
@@
For every such function $f$, we use the independence of $\xi^{\prime}$ and $(X,Y)$, along with the tower property and Jensen’s inequality to get
$\begin{split}\mathbb{E}[(Y-(P_{s}f)(\beta\circ X))^{2}]&=\mathbb{E}[(Y-\mathbb{E}[f(\beta\circ X+\sqrt{s}\,\beta\circ\xi)|X])^{2}]\\ &\leq\mathbb{E}[(Y-f(\beta\circ X+\sqrt{s}\,\beta\circ\xi))^{2}].\end{split}$ (4.1)
Let $\beta^{(s)}:=\beta-sv$ with $v=-\Pi_{S_{*}^{c}}\beta$. Then $\beta^{(s)}=\beta+s\Pi_{S_{*}^{c}}\beta$. We expand vectors in $\mathbb{R}^{d}$ by grouping coordinates in $S_{*}$ and $S_{*}^{c}$:
$\begin{split}\beta\circ X+\sqrt{s}\,\beta\circ\xi&=(\beta_{S_{*}}\circ X_{S_{*}},\;\beta_{S_{*}^{c}}\circ X_{S_{*}^{c}}+\sqrt{s}\,\beta_{S_{*}^{c}}\circ\xi^{\prime})\\ \beta^{(s)}\circ X&=(\beta_{S_{*}}\circ X_{S_{*}},\;(1+s)\beta_{S_{*}^{c}}\circ X_{S_{*}^{c}})\end{split}.$
Since $\xi^{\prime}$ has the same Gaussian distribution as $X_{S_{*}^{c}}$ and since $\xi^{\prime}$ is independent of $X$ whereas $X_{S_{*}}$ and $X_{S_{*}^{c}}$ are independent, we conclude the following equality in distribution:
$\begin{split}(X_{S_{*}},\beta\circ X+\sqrt{s}\,\beta\circ\xi)&\stackrel{{\scriptstyle d}}{{=}}(X_{S_{*}},\beta^{(s)}\circ X).\end{split}$
As a result,
$\mathbb{E}[(\mathbb{E}[Y|X_{S_{*}}]-f(\beta\circ X+\sqrt{s}\,\beta\circ\xi))^{2}]=\mathbb{E}[(\mathbb{E}[Y|X_{S_{*}}]-f(\beta^{(s)}\circ X))^{2}].$
@@
$\begin{split}\mathbb{E}[(Y-f(\beta\circ X+\sqrt{s}\,\beta\circ\xi))^{2}]&=\mathbb{E}[(Y-f(\beta^{(s)}\circ X))^{2}].\end{split}$ (4.2)
@@
$\mathbb{E}[(Y-(P_{s}f)(\beta\circ X))^{2}]\leq\mathbb{E}[(Y-f(\beta^{(s)}\circ X))^{2}].$ (4.3)
@@
Combining (4.3) and (4.4), we have shown for every $f\in H$,
$\mathbb{E}[(Y-(P_{s}f)(\beta\circ X))^{2}]+\lambda\left\|{P_{s}f}\right\|_{H}^{2}\leq\mathbb{E}[(Y-f(\beta^{(s)}\circ X))^{2}]+\lambda\left\|{f}\right\|_{H}^{2}.$
Thus, taking infimum over $f\in H$ on both sides yields $\mathcal{J}(\beta,\lambda)\leq\mathcal{J}(\beta^{(s)},\lambda)$. This yields
$\mathrm{D}\mathcal{J}(\beta,\lambda)[v]=-\lim_{s\to 0^{+}}\frac{1}{s}(\mathcal{J}(\beta^{(s)},\lambda)-\mathcal{J}(\beta,\lambda))\leq 0$
@@
$\begin{split}\liminf_{s\to 0^{+}}\frac{1}{s}(\left\|{f}\right\|_{H}^{2}-\left\|{P_{s}f}\right\|_{H}^{2})&\geq\liminf_{s\to 0^{+}}\int\frac{|\widehat{f}|(\omega)^{2}}{k(\omega)}(1-|\widehat{\pi}_{s}(-\omega_{S_{*}^{c}})|^{2})d\omega\\ &\geq 2\pi^{2}\cdot\int\frac{|\widehat{f}|(\omega)^{2}}{k(\omega)}\cdot\|(\beta\circ\omega)_{S_{*}^{c}}\|^{2}_{\Sigma}~d\omega\geq 0.\end{split}$
The quantity is strictly positive unless $f$ is zero. Since
$\begin{split}\mathcal{J}(\beta^{(s)},\lambda)=\mathcal{I}(\beta^{(s)},f_{\beta^{(s)},\lambda},\lambda)&=\mathbb{E}[(Y-f_{\beta^{(s)},\lambda}(\beta^{(s)}\circ X))^{2}]+\lambda\left\|{f_{\beta^{(s)},\lambda}}\right\|_{H}^{2}\\ \mathcal{J}(\beta,\lambda)\leq\mathcal{I}(\beta,P_{s}f_{\beta^{(s)},\lambda},\lambda)&=\mathbb{E}[(Y-(P_{s}f_{\beta^{(s)},\lambda})(\beta\circ X))^{2}]+\lambda\left\|{P_{s}f_{\beta^{(s)},\lambda}}\right\|_{H}^{2}\end{split}$

### v2

@@ Abstract
…..We study a compositional variant of kernel ridge regression in which the predictor is applied to a coordinate-wise reweighting of the inputs. Formulated as a variational problem, this model provides a tractable setting for studying feature learning in compositional architectures. From the perspective of variable selection, we show how relevant variables are recovered while noise variables are eliminated. We prove that both global minimizers and stationary points discard noise coordinates when the noise variables are Gaussian distributed. A central finding is that $\ell_{1}$-type kernels, such as the Laplace kernel, succeed in recovering features contributing to nonlinear effects at stationary points, whereas Gaussian kernels recover only linear ones.
@@ 1 Introduction
Our starting point is the classical kernel ridge regression (KRR) framework, which aims to learn a nonlinear function $f$ that predicts the response $Y\in\mathbb{R}$ from $X\in\mathbb{R}^{d}$, where $(X,Y)\sim\mathbb{P}$ [2]. The function $f$ is chosen from a reproducing kernel Hilbert space $H$ (e.g., Sobelev space). In KRR, the learning problem is formulated as the solution to the variational problem
@@ 1.2 Related Work
    Another recent line of work centers on the so-called AGOP (Average Gradient Outer Product) method [12]. This line of work also adopts the parameterization $f(\beta\circ X)$ (and $f(UX)$) within kernel ridge regression and can be viewed as operating on the same objective $\mathcal{J}$. AGOP relies on an iterative update scheme $\beta^{(t+1)}={\rm AGOP}(\beta^{(t)})$ (and $U^{(t+1)}={\rm AGOP}(U^{(t)})$ resp.) designed to extract a low-dimensional set of relevant features, and has been studied both theoretically—with generalization error bounds that outperform that of standard kernel methods—and empirically, with remarkable success on tasks such as tabular dataset predictions [12, 16]. Our study takes a complementary perspective. Instead of analyzing a specific iterative scheme, we focus on the optimization landscape of $\mathcal{J}$ itself and establish population-level guarantees on feature recovery for its global minimizers and stationary points. These results hold under more relaxed distributional assumptions and highlight how the geometry of $\mathcal{J}$ shapes variable selection independently of any algorithmic choices. A central insight of this analysis is that the choice of kernel is crucial: whereas existing guarantees primarily focus on smooth kernels, we show that nonsmooth $\ell_{1}$-type kernels such as the Laplace kernel enables the recovery of nonlinear main effects that Gaussian kernels miss. Interestingly, a recent algorithmic development on AGOP—namely xRFM [17], a state-of-the-art method for tabular datasets—demonstrates empirically that adopting the Laplace kernel improves generalization over other kernels (e.g., Gaussian), aligning with the theoretical insights developed in this paper.
@@ 1.4 Notation
Let $x\in\mathbb{R}^{d}$ be a vector. We write $x=(x_{1},\ldots,x_{d})$, where $x_{i}$ denotes the $i$-th coordinate of $x$, and define its support as $\mathop{\rm supp}(x):=\{i\in\{1,\ldots,d\}:x_{i}\neq 0\}$. For a subset $S\subset\{1,\ldots,d\}$, we write $x_{S}\in\mathbb{R}^{|S|}$ for the subvector of $x$ containing coordinates $\{x_{i}:i\in S\}$. We use $\|\cdot\|_{p}$ to denote the $\ell_{p}$ norm on $\mathbb{R}^{d}$. We use $e_{i}\in\mathbb{R}^{d}$ to denote the $i$-th standard basis vector. We use $\bar{z}$ to denote the complex conjugate of a complex number $z\in\mathbb{C}$. For complex-valued random variables $Z_{1},Z_{2}$, we define the covariance as ${\rm Cov}(Z_{1},Z_{2}):=\mathbb{E}[Z_{1}\bar{Z_{2}}]-\mathbb{E}[Z_{1}]\mathbb{E}[\bar{Z_{2}}]$, and write ${\rm Cov}(Z_{1}):={\rm Cov}(Z_{1},Z_{1})$.
@@ 2.1 Translation Invariant RKHS
Let $H$ be a Hilbert space of real-valued continuous function on $\mathbb{R}^{d}$ associated to a translation-invariant kernel. The space $H$ is defined in terms of the Fourier transform by
$H:=\left\{f:f\in L_{2}(\mathbb{R}^{d}),\left\|{f}\right\|_{H}^{2}=\int_{\mathbb{R}^{d}}\frac{|\hat{f}|^{2}(\omega)}{k(\omega)}d\omega<\infty\right\}$
@@
Lemma 1 (Continuous embedding of $H$ in $C(\mathbb{R}^{d})$).
For any $f\in H$, we have $\left\|{f}\right\|_{L_{\infty}}\leq\left\|{f}\right\|_{H}$, and $f$ is continuous, and $\lim_{x\to\infty}f(x)=0$.
@@
Since $\hat{f}\in L_{1}$, the continuity of $f$ follows from Lebesgue’s dominated convergence theorem. By the Riemann-Lebesgue lemma, we get $\lim_{x\to\infty}f(x)=0$. ∎
@@
Suppose this is not true. There is nonzero $f\in L_{2}(\mu)$ such that $\int fhd\mu=0$ for all $h\in H$. In particular, taking $h(\cdot)=K(\cdot-z)\in H$, we obtain $\int f(x)K(x-z)\mu(dx)=0$ for all $z\in\mathbb{R}^{d}$. Using the Fourier inversion formula, $K(x-z)=\int k(\omega)e^{2\pi i\langle x-z,\omega\rangle}d\omega$, we get
$\begin{split}\iint f(x)k(\omega)e^{2\pi i\langle x-z,\omega\rangle}d\omega\mu(dx)=0~~~\forall z\in\mathbb{R}^{d}.\end{split}$
Define the Fourier transform of $f$ under $\mu$ as $F(\omega):=\int f(x)e^{-2\pi i\langle x,\omega\rangle}\mu(dx)$. Then the above equation becomes
$\int F(-\omega)k(\omega)e^{-2\pi i\langle z,\omega\rangle}d\omega=0~~~\forall z\in\mathbb{R}^{d}.$
That is, the Fourier transform of $F(-\omega)k(\omega)$ is identically zero. Since $k(\omega)>0$, this implies $F(\omega)=0$ almost everywhere $\omega$, so $f=0$ in $L_{2}(\mu)$—a contradiction. ∎
@@ 2.3 Continuity in Parameter
A fundamental property is that the minimizer $f$ and the objective $\mathcal{J}$ depend continuously on the parameter $\beta$. This fact essentially follows from the reinterpretation of the $f(UX)$ model in related work [14]; for completeness, we include a self-contained proof in this manuscript.
@@
Since $f_{\beta,\lambda}\in H$ is continuous and uniformly bounded in $\mathbb{R}^{d}$ by Lemma 1, we get $\mathcal{I}(f_{\beta,\lambda},\beta^{\prime},\lambda)\to\mathcal{I}(f_{\beta,\lambda},\beta,\lambda)=\mathcal{J}(\beta,\lambda)$ as $\beta^{\prime}\to\beta$. Also, by Lemma 8, we have $\mathcal{J}(\beta^{\prime},\lambda)\to\mathcal{J}(\beta,\lambda)$. This implies
@@ 2.4 First Variation
Related formulas have appeared in the author’s earlier work on $f(UX)$ model [14, Proposition 4.4], though under stronger smoothness assumptions, i.e., full Fréchet differentiability on the map $\beta\mapsto K_{\beta}(x,x^{\prime})$ for fixed $x,x^{\prime}$, and in more specialized settings where $K$ is assumed to be a radial basis kernel.
@@ 2.4.1 Examples
Theorem 1 and the above examples illustrate that, in many scenarios of interest, the directional derivative $\mathrm{D}\mathcal{J}(\beta,\lambda)[v]$ exists for all $\beta\in\mathbb{R}^{d}$ and all $v\in\mathbb{R}^{d}$. Accordingly, throughout the paper we assume this derivative exists for all $\beta$ and $v$.
@@ 3.1 Existence
Continuity of $X$ serves as a sufficient condition for $\mathcal{J}$ to admit a finite minimizer. In the presence of discrete variables, it may happen that $\limsup_{\beta\to\infty}\mathcal{J}(\beta,\lambda)<\mathbb{E}[Y^{2}]$ (see [14, Section 5]). However, we are unaware of a situation where there is a global minimizer at infinity—a question we leave open.
@@ 3.2.1 Core Sufficient Feature Set
To study feature learning rigorously, we focus on settings where a well-defined target structure exists. Accordingly, we assume throughout:
@@
Under Assumption 3.1, the set $S_{*}$ provides a meaningful statistical target: the minimal subset of input features needed to predict $Y$. The problem of variable selection concerns whether minimizing $\mathcal{J}$ successfully recovers $S_{*}$. Understanding when and why this recovery succeeds is key to revealing how the compositional model $f(\beta\circ X)$ learns structure from data.
@@ 3.2.2 Irrelevant Feature Elimination
We first study one aspect of feature learning: when does the global minimizer $\beta_{*}$ eliminate irrelevant features—that is, assign zero weight to coordinates outside the core sufficient feature set $S_{*}$? Such sparsity reflects the model’s ability to suppress irrelevant inputs automatically.
@@
This assumption posits that the relevant features in $X$ for predicting YY—those in $S_{*}$—are independent of the rest. Such independence avoids potential statistical confounding, making it clear in principle to distinguish relevant features based solely on their predictive contribution.
@@
Assume $\mathbb{E}[Y|X]\neq 0$. Let $\mathbf{1}\in\mathbb{R}^{d}$ denote the all-ones vector. We show that $\mathcal{J}(\mathbf{1},\lambda)<\mathbb{E}[Y^{2}]$. Suppose instead that $\mathcal{J}(\mathbf{1},\lambda)=\mathbb{E}[Y^{2}]$. Then the minimizer $f_{\mathbf{1},\lambda}$ must be identically zero, which by Lemma 4 implies $\mathbb{E}[Yg(X)]=0$ for all $g\in H$. Since $H$ is dense in $L_{2}(\mathbb{P}_{X})$ by Lemma 2, it follows that $\mathbb{E}[Yg(X)]=0$ for all $g\in L_{2}(\mathbb{P}_{X})$, and hence $\mathbb{E}[Y|X]=0$ almost surely—contradicting our assumption $\mathbb{E}[Y|X]\neq 0$. We conclude that $\mathcal{J}(\mathbf{1},\lambda)<\mathbb{E}[Y^{2}]$, and thus $\inf_{\beta\in\mathbb{R}^{d}}\mathcal{J}(\beta,\lambda)<\mathbb{E}[Y^{2}]$.
@@ 3.2.3 Relevant Feature Recovery
On the other hand, $H$ is dense in $L_{2}(\mathbb{P}_{X})$ by Lemma 2, where $\mathbb{P}_{X}$ denotes the distribution of $X$. Thus, the regression function $x\mapsto\mathbb{E}[Y|X=x]$ can be approximated arbitrarily well in $\mathcal{L}_{2}(\mathbb{P}_{X})$ by functions in $H$. In particular, there exists $f_{\varepsilon}\in H$ such that
@@ 4 Stationary Points and Irrelevant Feature Elimination
Unlike global minimizers, stationary points are harder to characterize, and require more delicate analytical techniques. To proceed, we divide the analysis into two parts. This section addresses the following question:
@@
Since the problem is translation-invariant, we can without loss of generality assume $\mathbb{E}[X]=0$. Let $\xi\in\mathbb{R}^{d}$ be a Gaussian random vector independent of $(X,Y)$ with
@@
Here and throughout, we use $\stackrel{{\scriptstyle d}}{{=}}$ to denote equality in distribution.
Fix $\beta\in\mathbb{R}^{d}$. For $s\in[0,1]$, define the vector $\beta^{(s)}\in\mathbb{R}^{d}$ by
$\beta^{(s)}=(\beta_{S_{*}},\sqrt{1-s}\,\beta_{S_{*}^{c}}).$
Writing Hadamard products blockwise, we have
$\begin{split}\beta\circ X&=(\beta_{S_{*}}\circ X_{S_{*}},\;\beta_{S_{*}^{c}}\circ X_{S_{*}^{c}})\\ \beta^{(s)}\circ X+\sqrt{s}\,\beta\,\circ\,\xi&=(\beta_{S_{*}}\circ X_{S_{*}},\;\beta_{S_{*}^{c}}\circ(\sqrt{1-s}\,X_{S_{*}^{c}}+\sqrt{s}\,\xi^{\prime}))\end{split}.$
Since $\xi^{\prime}$ and $X_{S_{*}^{c}}$ are independent and identically Gaussian distributed, we get $X_{S_{*}^{c}}\stackrel{{\scriptstyle d}}{{=}}\sqrt{1-s}\,X_{S_{*}^{c}}+\sqrt{s}\,\xi^{\prime}$. Moreover, since $X_{S_{*}}$ is independent of both $X_{S_{*}^{c}}$ and $\xi^{\prime}$, it follows that
$(X_{S_{*}},\,\beta\circ X)\stackrel{{\scriptstyle d}}{{=}}(X_{S_{*}},\,\beta^{(s)}\circ X+\sqrt{s}\,\beta\,\circ\,\xi).$ (4.1)
Let $\pi_{s}$ denote the probability measure of the random vector $\sqrt{s}\,\beta_{S_{*}^{c}}\,\circ\,\xi^{\prime}\in\mathbb{R}^{|S_{*}^{c}|}$. For every $f\in H$ and $s>0$, we define $P_{s}$ as the convolution type operator:
@@
For every such function $f$, we use the independence of $\xi$ and $(X,Y)$ to get
$Y-(P_{s}f)(\beta^{(s)}\circ X)=\mathbb{E}[Y-f(\beta^{(s)}\circ X+\sqrt{s}\,\beta\circ\xi)|X,Y].$
By Jensen’s inequality and tower property, we further get
$\begin{split}\mathbb{E}[(Y-(P_{s}f)(\beta^{(s)}\circ X))^{2}]&=\mathbb{E}[(\mathbb{E}[Y-f(\beta^{(s)}\circ X+\sqrt{s}\,\beta\circ\,\xi)|X,Y])^{2}]\\ &\leq\mathbb{E}[\mathbb{E}[(Y-f(\beta^{(s)}\circ X+\sqrt{s}\,\beta\circ\,\xi))^{2}|X,Y]]\\ &=\mathbb{E}[(Y-f(\beta^{(s)}\circ X+\sqrt{s}\,\beta\circ\,\xi))^{2}].\end{split}$ (4.2)
We evaluate the right-hand-side of the equation. By the equality in distribution in (4.1),
$\mathbb{E}[(\mathbb{E}[Y|X_{S_{*}}]-f(\beta^{(s)}\circ X+\sqrt{s}\,\beta\circ\xi))^{2}]=\mathbb{E}[(\mathbb{E}[Y|X_{S_{*}}]-f(\beta\circ X))^{2}].$
@@
$\begin{split}\mathbb{E}[(Y-f(\beta^{(s)}\circ X+\sqrt{s}\,\beta\circ\xi))^{2}]&=\mathbb{E}[(Y-f(\beta\circ X))^{2}].\end{split}$ (4.3)
@@
$\mathbb{E}[(Y-(P_{s}f)(\beta^{(s)}\circ X))^{2}]\leq\mathbb{E}[(Y-f(\beta\circ X))^{2}].$ (4.4)
@@
Hence, $P_{s}f\in H$. Combining (4.4) and (4.5), we have shown for every $f\in H$,
$\mathbb{E}[(Y-(P_{s}f)(\beta^{(s)}\circ X))^{2}]+\lambda\left\|{P_{s}f}\right\|_{H}^{2}\leq\mathbb{E}[(Y-f(\beta\circ X))^{2}]+\lambda\left\|{f}\right\|_{H}^{2}.$
Thus, taking infimum over $f\in H$ on both sides yields
$\mathcal{J}(\beta^{(s)},\lambda)\leq\mathcal{J}(\beta,\lambda).$ (4.6)
Let $v:=-\Pi_{S_{*}^{c}}\beta\in\mathbb{R}^{d}$. Since $\beta^{(s)}=\beta+\frac{1}{2}sv+o(|s|)$ as $s\to 0^{+}$, this yields
$\mathrm{D}\mathcal{J}(\beta,\lambda)[v]=\lim_{s\to 0^{+}}\frac{2}{s}(\mathcal{J}(\beta^{(s)},\lambda)-\mathcal{J}(\beta,\lambda))\leq 0$
@@
$\begin{split}\liminf_{s\to 0^{+}}\frac{1}{s}(\left\|{f}\right\|_{H}^{2}-\left\|{P_{s}f}\right\|_{H}^{2})&\geq\liminf_{s\to 0^{+}}\int\frac{|\widehat{f}|(\omega)^{2}}{k(\omega)}(1-|\widehat{\pi}_{s}(-\omega_{S_{*}^{c}})|^{2})d\omega\\ &\geq 2\pi^{2}\cdot\int\frac{|\widehat{f}|(\omega)^{2}}{k(\omega)}\cdot\|(\beta\circ\omega)_{S_{*}^{c}}\|^{2}_{\Sigma}~d\omega.\end{split}$
Since
$\begin{split}\mathcal{J}(\beta^{(s)},\lambda)\leq\mathcal{I}(\beta^{(s)},P_{s}f_{\beta,\lambda},\lambda)&=\mathbb{E}[(Y-(P_{s}f_{\beta,\lambda})(\beta^{(s)}\circ X))^{2}]+\lambda\left\|{P_{s}f_{\beta,\lambda}}\right\|_{H}^{2}\\ \mathcal{J}(\beta,\lambda)=\mathcal{I}(\beta,f_{\beta,\lambda},\lambda)&=\mathbb{E}[(Y-f_{\beta,\lambda}(\beta\circ X))^{2}]+\lambda\left\|{f_{\beta,\lambda}}\right\|_{H}^{2}\end{split}$
and since by (4.4)
$\mathbb{E}[(Y-(P_{s}f_{\beta,\lambda})(\beta^{(s)}\circ X))^{2}]\leq\mathbb{E}[(Y-f_{\beta,\lambda}(\beta\circ X))^{2}]$